# Guaranteeing Semantic and Performance Determinism in Flexible GPU Sharing

### 16 Mar 2026 v1

@@ 1 Introduction
Artificial intelligence (AI) applications, such as chatbots [openai2023chatgpt, team2023gemini], code generation [jiang2024self, gu2023llm], and autonomous agents [zhao2024expel], have become the dominant consumers of cloud compute resources and rely heavily on powerful GPU accelerators [Gartner2025AI_IaaS, Hyperion2024CloudAIAI, li2025ecoserve]. These workloads exhibit highly dynamic resource demands, frequently alternating between compute-bound phases (e.g., prompt prefill) and memory-bound phases (e.g., autoregressive decoding). Current cloud abstractions fail to match this dynamism, primarily offering rigid, fixed-size instances to enforce isolation. This mismatch compels operators to over-provision for peak loads, resulting in significant resource waste during off-peak cycles. For instance, Microsoft reports that compute utilization drops to under 10% during the memory-bound decoding phase when serving Llama3-8B [llama3] on A100 GPUs [kamath2025pod]. Industry surveys indicate that 60%–70% of GPU capacity routinely remains idle across both training and inference workloads in production clusters, translating to billions of dollars in wasted capital expenditure annually. Attempts to mitigate this inefficiency via manual resource tuning prove impractical, as the process is labor-intensive and fails to scale [Mirantis, DevZero].
@@
GPU resource sharing offers a natural solution: by co-locating workloads, the capacity left idle during one task’s low-utilization phases can be reclaimed by co-resident jobs. Existing approaches can be broadly classified into two categories. Temporal sharing [wu2023transparent, xiao2018gandiva, xiao2020antman, duato2011enabling] time-slices the GPU, granting each job exclusive access for fixed quanta. While straightforward, this approach fails to reduce under-utilization during memory-bound phases and can cause severe latency spikes for latency-critical services [characterizing, gilman2021demystifying]. Spatial sharing [gslice, choi2022serving, han2024kace, dhakal2020spatial, zhang2025improving] permits concurrent execution by partitioning GPU resources (e.g., Streaming Multiprocessors (SMs)) across workloads. This approach is considered more promising for achieving high utilization, yet existing spatial-sharing systems remain impractical due to three main challenges:
@@
  • C1: Semantic Determinism—Inequivalence after sharing. Fine-grained spatial sharing approaches [gslice, coppock2025lithos] often modify kernel code, fundamentally altering execution semantics. They change either the effective batch size (distorting normalization statistics) or the computation order (violating floating-point associativity), causing numerical divergence from standalone execution [he2025nondeterminism].
@@
  • C2: Performance Determinsim—Unpredictable SLO violations. Co-locating services on shared GPUs inherently introduces resource contention. Compounding this, existing infrastructure [yu2022orca, patel2024splitwise] suffers from loss of semantic information: it remains oblivious to request priority. This uniform handling exacerbates interference, jeopardizing strict tail-latency Service Level Objectives (SLOs).
@@
  • C3: Transparency and Practicality—The engineering tax. Techniques like searching for fusible kernels [strati2024orion] or manual kernel rewriting [coppock2025lithos] require deep domain expertise and invasive code modifications, imposing an engineering tax. This not only creates maintenance burdens across rapidly evolving model architectures and underlying software stacks, but also amplifies the risk of violating semantic determinism (C1).
@@
This paper presents DetShare, a transparent GPU sharing system that simultaneously addresses all three challenges. Our key observation is that the modern GPU software-hardware stack already offers the necessary primitives for efficient spatial sharing without sacrificing semantic determinism: NVIDIA’s Multi-Process Service (MPS) [nvidia_mps_guide] enables spatial sharing with address space isolation, while Green Contexts [NVIDIA2025Green] serve as lightweight execution units amenable to fine-grained preemption.
@@ 2.1 AI Model Serving
AI models are now ubiquitous across modern applications, powering search, personalization, automation, diagnostics, and conversational agents. And model serving have become primary consumers of computing resources [Gartner2025AI_IaaS, Hyperion2024CloudAIAI, li2025ecoserve]. Serving AI models involves two broad phases, training produces a model and inference makes the model’s predictions available to various applications.
Training includes pre-training, referring to large-scale, often self-supervised training on massive data to build foundation models [devlin2019bert, bao2021beit, li2023blip]. These pre-trained foundation models then serve as general-purpose backbones for downstream applications in natural language understanding, conversational agents, summarization, image representation learning, and multimodal reasoning. Post-training further adapts foundation models to domain-specific tasks through fine-tuning [howard2018universal, ding2023parameter], instruction tuning [wei2021finetuned, peng2023instruction, liu2023visual] or personalization [zhuang2024hydra, zhang2019deep]. In addition, model compression [han2016deep, hinton2015distilling] and quantization [banner2019post] techniques are applied to reduce memory footprint and improve inference latency. Another kind of training is online or continual training [parisi2019continual], incrementally updating the deployed models using fresh or user-generated data.
Inference entails the execution of trained AI models to generate predictions for online or batch tasks. Within this domain, LLM serving has emerged as a dominant workload, underpinning a diverse array of critical applications ranging from conversational agents [openai2023chatgpt, team2023gemini] and neural code synthesis [jiang2024self, gu2023llm] to Retrieval-Augmented Generation (RAG) pipelines [lewis2020retrieval, gao2023retrieval] and autonomous agentic systems [zhao2024expel]. Typical LLM inference can be conceptually divided into a prefill phase, where the input context is encoded and key-value caches are initialized, and a decode phase, where tokens are generated sequentially, often with beam search or sampling strategies. These phases exhibit distinct resource utilization patterns: prefill is primarily compute-bound, characterized by high arithmetic intensity due to parallel processing of input tokens. In contrast, decode is memory-bound, constrained by memory bandwidth due to the low arithmetic intensity of auto-regressive token generation and the overhead of KV cache loading [agrawal2024taming, zhu2025nanoflow, kamath2025pod].
@@ 2.2 GPU Sharing
Temporal sharing [wu2023transparent, xiao2018gandiva, xiao2020antman, duato2011enabling] allows one to switch from one application to another. This offers strong isolation, while context switching introduces heavy overhead and since only one context is activated, overall GPU utilization can be low when the workload is small [ausavarungnirun2018mask, bai2020pipeswitch].
Spatial sharing [wu2023transparent, xiao2018gandiva, xiao2020antman, duato2011enabling] executes multiple workloads concurrently across partitioned GPU resources, spanning from fine-grained software multiplexing like CUDA Streams to coarse-grained hardware isolation. While CUDA Streams [choi2022serving, lim2021zico, pavlidakis2022arax, wang2021wavelet] enable intra-application parallelism by issuing operations into distinct hardware work queues, they operate within a single CUDA context. Although this eliminates context switching overhead, the resulting unified address space enforces fate sharing: a fault in one stream crashes the entire context, offering zero fault isolation.
To provide stronger multi-tenant isolation, vendors offer dedicated spatial partitioning mechanisms like NVIDIA’s Multi-Instance GPU (MIG) and Multi-Process Service (MPS). MIG [nvidia_mig_guide] achieves strict hardware-level isolation by dividing a single physical GPU into fully isolated instances, each with dedicated compute and memory bandwidth. However, this isolation comes at a severe cost: MIG relies on heavyweight reconfiguration mechanisms that necessitate device resets, incurring latencies on the order of hundreds of milliseconds [strati2024orion, li2022miso]. Furthermore, its coarse-grained partitioning enforces fixed resource slice sizes, inevitably leading to significant resource fragmentation. In contrast to hardware-centric partitioning, MPS [nvidia_mps_guide] operates at the software level, facilitating process-level spatial multiplexing via a client-server architecture. By multiplexing command streams from multiple independent clients into a single shared hardware context, MPS eliminates context-switching overheads while successfully maintaining separate virtual address spaces for each process. This design makes MPS particularly beneficial for multi-tenant inference workloads where clients can safely co-exist within a context, effectively boosting overall utilization by absorbing leftover parallelism without the rigid reconfiguration penalties of MIG.
@@ 2.3 The Sources of Semantic Non-Determinism
At the lowest level, the root cause of numerical divergence is the IEEE 754 standard for floating-point arithmetic [goldberg1991every]. Due to finite precision (especially in truncated formats like FP16 and BF16), floating-point addition is non-associative: $(A+B)+C\neq A+(B+C)$. Consequently, any variation in the sequence of arithmetic reductions will yield divergent rounding errors.
@@
To maximize throughput, modern LLM serving engines dynamically adapt execution paths, which inherently introduces non-determinism. Techniques such as continuous batching [yu2022orca] cause effective batch sizes to fluctuate heavily. Concurrently, high-performance computational kernels (e.g., cuBLAS) leverage dynamic tile sizing and Split-K reductions for matrix multiplications [he2025nondeterminism, nvidia_cublas]. By continually reshaping the underlying reduction trees, these performance-oriented optimizations alter the sequence of floating-point operations, directly exacerbating numerical divergence.
@@
This numerical non-determinism extends beyond mere theoretical artifacts; it severely impacts advanced training paradigms such as Reinforcement Learning from Human Feedback (RLHF). During RL fine-tuning, identical prompts must consistently yield identical token trajectories. Micro-divergences introduced during inference trigger a policy shift [petrenko2026entropypreservingreinforcementlearning]—the model’s behavior during online rollouts deviates from the policy evaluated during offline training. This mismatch effectively degrades online RL algorithms into suboptimal offline RL, substantially stalling convergence.
Recognizing this, recent serving frameworks (e.g., SGLang [sglang2024, sglang_deterministic_blog]) and libraries [RepDL] have introduced strict “deterministic modes”. However, these application-level guarantees implicitly assume dedicated, unfragmented hardware. As we discuss in Section 3.1, when these workloads are subjected to spatial multiplexing and sharing mechanisms at the infrastructure layer, the semantic determinism meticulously preserved by the application is once again destroyed.
@@ 3.1 Challenge-#1: Semantic Determinism
The bedrock of production AI systems is semantic integrity: a model validated in an offline environment must behave identically when deployed online. The core requirement for GPU resource sharing, therefore, is behavioral equivalence—preserving each workload’s functional behavior during shared execution compared to its standalone counterpart. This property is paramount for training-inference consistency [liu-li-2025-rl-collapse]: models validated on dedicated hardware must yield identical predictions in inference scenarios. Any numerical divergence—however minute—can invalidate offline accuracy metrics. While temporal sharing naturally preserves arithmetic determinism by granting exclusive access to full resources, spatial sharing introduces complex challenges.
To maximize utilization, recent intrusive methods (e.g., LithOS [coppock2025lithos]) employ kernel atomization—dynamically partitioning parallelizable dimensions to fit available resource fragments. We argue that such techniques fundamentally violate the abstraction boundary between logical computation and physical resource availability. Specifically, standard GPU kernels operate on a contract where the result is a function of input data ($D$) and model state ($S$): $Result=F(D,S)$. However, by dynamically reshaping kernels to fit hardware fragments, atomization implicitly leaks instantaneous resource constraints ($R(t)$) into execution semantics, effectively altering the function to $Result^{\prime}=F(D,S,R(t))$. Since $R(t)$ is stochastic in a multi-tenant cluster, the computation becomes non-deterministic. We categorize this divergence into two distinct levels:
@@
  • Statistical Divergence (Batch Sensitivity): Techniques that partition resources by splitting batches effectively alter the statistical population used for operations like Batch Normalization [ioffe2015batch]. As demonstrated in Figure 1, reducing the effective batch size to fit fragmented resources distorts mean and variance statistics, leading to significant loss fluctuations and degraded convergence.
  • Numerical Divergence (FP Non-associativity): A more insidious issue arises in reduction-heavy kernels (e.g., Softmax, LayerNorm). Due to the non-associativity of floating-point arithmetic [goldberg1991every], the order of summation determines rounding error accumulation. As shown in Figure 2, we simulate resource fragmentation by splitting a large tensor summation into varying grid sizes. The results demonstrate that merely altering the reduction tree structure causes the absolute error (vs. standalone execution) to fluctuate significantly. This divergence is present in both FP16 and BF16, with BF16 showing an order of magnitude higher error, confirming that fragmentation breaks bitwise determinism.
@@ 3.2 Challenge-#2: Performance Determinism
Meeting strict tail-latency SLOs is a prerequisite for ensuring satisfactory user experience in interactive deep learning workloads. However, achieving performance determinism—specifically, keeping P99 latency stable—is notoriously difficult in multi-tenant environments. While temporal sharing is the incumbent solution, it is fundamentally ill-suited for strict latency constraints. The primary limitation is that time-slicing enforces serialization: it introduces unavoidable queuing delays for applications not currently scheduled on the GPU [clock, xiao2020antman]. As contention increases, these accumulated wait times frequently cause tail latencies to exceed strict deadlines.
@@
  • Micro-architectural Contention: Spatial co-location implicitly multiplexes shared resources that are not explicitly partitioned, most notably Device Memory Bandwidth and Last-Level Cache capacity. As characterized in prior studies [characterizing], a latency-sensitive inference job can suffer non-deterministic latency spikes (up to $3\times$ [kamath2025pod]) when co-located with a bandwidth-heavy training job.
  • Scheduler Obliviousness: Commodity GPU hardware schedulers are throughput-oriented and oblivious to high-level SLO constraints. Without explicit intervention, a high-priority kernel may suffer from head-of-line blocking behind a massive, non-preemptible background kernel in the hardware command queue [xsched], rendering logical priority ineffective.

### 17 Mar 2026 v2

@@ 1 Introduction
Artificial intelligence (AI) applications, such as chatbots [56, 69], code generation [40, 26], and autonomous agents [84], have become the dominant consumers of cloud compute resources and rely heavily on powerful GPU accelerators [20, 65, 47]. These workloads exhibit highly dynamic resource demands, frequently alternating between compute-bound phases (e.g., prompt prefill) and memory-bound phases (e.g., autoregressive decoding). Current cloud abstractions fail to match this dynamism, primarily offering rigid, fixed-size instances to enforce isolation. This mismatch compels operators to over-provision for peak loads, resulting in significant resource waste during off-peak cycles. For instance, Microsoft reports that compute utilization drops to under 10% during the memory-bound decoding phase when serving Llama3-8B [18] on A100 GPUs [43]. Industry surveys indicate that 60%–70% of GPU capacity routinely remains idle across both training and inference workloads in production clusters, translating to billions of dollars in wasted capital expenditure annually. Attempts to mitigate this inefficiency via manual resource tuning prove impractical, as the process is labor-intensive and fails to scale [39, 74].
@@
GPU resource sharing offers a natural solution: by co-locating workloads, the capacity left idle during one task’s low-utilization phases can be reclaimed by co-resident jobs. Existing approaches can be broadly classified into two categories. Temporal sharing [75, 76, 77, 17] time-slices the GPU, granting each job exclusive access for fixed quanta. While straightforward, this approach fails to reduce under-utilization during memory-bound phases and can cause severe latency spikes for latency-critical services [22, 21]. Spatial sharing [15, 8, 30, 14, 83] permits concurrent execution by partitioning GPU resources (e.g., Streaming Multiprocessors (SMs)) across workloads. This approach is considered more promising for achieving high utilization, yet existing spatial-sharing systems remain impractical due to three main challenges:
@@
  • C1: Semantic Determinism—Inequivalence after sharing. Fine-grained spatial sharing approaches [15, 9] often modify kernel code, fundamentally altering execution semantics. They change either the effective batch size (distorting normalization statistics) or the computation order (violating floating-point associativity), causing numerical divergence from standalone execution [32].
@@
  • C2: Performance Determinsim—Unpredictable SLO violations. Co-locating services on shared GPUs inherently introduces resource contention. Compounding this, existing infrastructure [79, 58] suffers from loss of semantic information: it remains oblivious to request priority. This uniform handling exacerbates interference, jeopardizing strict tail-latency Service Level Objectives (SLOs).
@@
  • C3: Transparency and Practicality—The engineering tax. Techniques like searching for fusible kernels [67] or manual kernel rewriting [9] require deep domain expertise and invasive code modifications, imposing an engineering tax. This not only creates maintenance burdens across rapidly evolving model architectures and underlying software stacks, but also amplifies the risk of violating semantic determinism (C1).
@@
This paper presents DetShare, a transparent GPU sharing system that simultaneously addresses all three challenges. Our key observation is that the modern GPU software-hardware stack already offers the necessary primitives for efficient spatial sharing without sacrificing semantic determinism: NVIDIA’s Multi-Process Service (MPS) [10] enables spatial sharing with address space isolation, while Green Contexts [55] serve as lightweight execution units amenable to fine-grained preemption.
@@ 2.1 AI Model Serving
AI models are now ubiquitous across modern applications, powering search, personalization, automation, diagnostics, and conversational agents. And model serving have become primary consumers of computing resources [20, 65, 47]. Serving AI models involves two broad phases, training produces a model and inference makes the model’s predictions available to various applications.
Training includes pre-training, referring to large-scale, often self-supervised training on massive data to build foundation models [13, 6, 46]. These pre-trained foundation models then serve as general-purpose backbones for downstream applications in natural language understanding, conversational agents, summarization, image representation learning, and multimodal reasoning. Post-training further adapts foundation models to domain-specific tasks through fine-tuning [35, 16], instruction tuning [73, 60, 49] or personalization [87, 82]. In addition, model compression [31, 34] and quantization [5] techniques are applied to reduce memory footprint and improve inference latency. Another kind of training is online or continual training [57], incrementally updating the deployed models using fresh or user-generated data.
Inference entails the execution of trained AI models to generate predictions for online or batch tasks. Within this domain, LLM serving has emerged as a dominant workload, underpinning a diverse array of critical applications ranging from conversational agents [56, 69] and neural code synthesis [40, 26] to Retrieval-Augmented Generation (RAG) pipelines [44, 19] and autonomous agentic systems [84]. Typical LLM inference can be conceptually divided into a prefill phase, where the input context is encoded and key-value caches are initialized, and a decode phase, where tokens are generated sequentially, often with beam search or sampling strategies. These phases exhibit distinct resource utilization patterns: prefill is primarily compute-bound, characterized by high arithmetic intensity due to parallel processing of input tokens. In contrast, decode is memory-bound, constrained by memory bandwidth due to the low arithmetic intensity of auto-regressive token generation and the overhead of KV cache loading [1, 86, 43].
@@ 2.2 GPU Sharing
Temporal sharing [75, 76, 77, 17] allows one to switch from one application to another. This offers strong isolation, while context switching introduces heavy overhead and since only one context is activated, overall GPU utilization can be low when the workload is small [2, 4].
Spatial sharing [75, 76, 77, 17] executes multiple workloads concurrently across partitioned GPU resources, spanning from fine-grained software multiplexing like CUDA Streams to coarse-grained hardware isolation. While CUDA Streams [8, 48, 59, 71] enable intra-application parallelism by issuing operations into distinct hardware work queues, they operate within a single CUDA context. Although this eliminates context switching overhead, the resulting unified address space enforces fate sharing: a fault in one stream crashes the entire context, offering zero fault isolation.
To provide stronger multi-tenant isolation, vendors offer dedicated spatial partitioning mechanisms like NVIDIA’s Multi-Instance GPU (MIG) and Multi-Process Service (MPS). MIG [11] achieves strict hardware-level isolation by dividing a single physical GPU into fully isolated instances, each with dedicated compute and memory bandwidth. However, this isolation comes at a severe cost: MIG relies on heavyweight reconfiguration mechanisms that necessitate device resets, incurring latencies on the order of hundreds of milliseconds [67, 45]. Furthermore, its coarse-grained partitioning enforces fixed resource slice sizes, inevitably leading to significant resource fragmentation. In contrast to hardware-centric partitioning, MPS [10] operates at the software level, facilitating process-level spatial multiplexing via a client-server architecture. By multiplexing command streams from multiple independent clients into a single shared hardware context, MPS eliminates context-switching overheads while successfully maintaining separate virtual address spaces for each process. This design makes MPS particularly beneficial for multi-tenant inference workloads where clients can safely co-exist within a context, effectively boosting overall utilization by absorbing leftover parallelism without the rigid reconfiguration penalties of MIG.
@@ 2.3 The Sources of Semantic Non-Determinism
At the lowest level, the root cause of numerical divergence is the IEEE 754 standard for floating-point arithmetic [25]. Due to finite precision (especially in truncated formats like FP16 and BF16), floating-point addition is non-associative: $(A+B)+C\neq A+(B+C)$. Consequently, any variation in the sequence of arithmetic reductions will yield divergent rounding errors.
@@
To maximize throughput, modern LLM serving engines dynamically adapt execution paths, which inherently introduces non-determinism. Techniques such as continuous batching [79] cause effective batch sizes to fluctuate heavily. Concurrently, high-performance computational kernels (e.g., cuBLAS) leverage dynamic tile sizing and Split-K reductions for matrix multiplications [32, 54]. By continually reshaping the underlying reduction trees, these performance-oriented optimizations alter the sequence of floating-point operations, directly exacerbating numerical divergence.
@@
This numerical non-determinism extends beyond mere theoretical artifacts; it severely impacts advanced training paradigms such as Reinforcement Learning from Human Feedback (RLHF). During RL fine-tuning, identical prompts must consistently yield identical token trajectories. Micro-divergences introduced during inference trigger a policy shift [61]—the model’s behavior during online rollouts deviates from the policy evaluated during offline training. This mismatch effectively degrades online RL algorithms into suboptimal offline RL, substantially stalling convergence.
Recognizing this, recent serving frameworks (e.g., SGLang [85, 51]) and libraries [78] have introduced strict “deterministic modes”. However, these application-level guarantees implicitly assume dedicated, unfragmented hardware. As we discuss in Section 3.1, when these workloads are subjected to spatial multiplexing and sharing mechanisms at the infrastructure layer, the semantic determinism meticulously preserved by the application is once again destroyed.
@@ 3.1 Challenge-#1: Semantic Determinism
The bedrock of production AI systems is semantic integrity: a model validated in an offline environment must behave identically when deployed online. The core requirement for GPU resource sharing, therefore, is behavioral equivalence—preserving each workload’s functional behavior during shared execution compared to its standalone counterpart. This property is paramount for training-inference consistency [50]: models validated on dedicated hardware must yield identical predictions in inference scenarios. Any numerical divergence—however minute—can invalidate offline accuracy metrics. While temporal sharing naturally preserves arithmetic determinism by granting exclusive access to full resources, spatial sharing introduces complex challenges.
To maximize utilization, recent intrusive methods (e.g., LithOS [9]) employ kernel atomization—dynamically partitioning parallelizable dimensions to fit available resource fragments. We argue that such techniques fundamentally violate the abstraction boundary between logical computation and physical resource availability. Specifically, standard GPU kernels operate on a contract where the result is a function of input data ($D$) and model state ($S$): $Result=F(D,S)$. However, by dynamically reshaping kernels to fit hardware fragments, atomization implicitly leaks instantaneous resource constraints ($R(t)$) into execution semantics, effectively altering the function to $Result^{\prime}=F(D,S,R(t))$. Since $R(t)$ is stochastic in a multi-tenant cluster, the computation becomes non-deterministic. We categorize this divergence into two distinct levels:
@@
  • Statistical Divergence (Batch Sensitivity): Techniques that partition resources by splitting batches effectively alter the statistical population used for operations like Batch Normalization [38]. As demonstrated in Figure 1, reducing the effective batch size to fit fragmented resources distorts mean and variance statistics, leading to significant loss fluctuations and degraded convergence.
  • Numerical Divergence (FP Non-associativity): A more insidious issue arises in reduction-heavy kernels (e.g., Softmax, LayerNorm). Due to the non-associativity of floating-point arithmetic [25], the order of summation determines rounding error accumulation. As shown in Figure 2, we simulate resource fragmentation by splitting a large tensor summation into varying grid sizes. The results demonstrate that merely altering the reduction tree structure causes the absolute error (vs. standalone execution) to fluctuate significantly. This divergence is present in both FP16 and BF16, with BF16 showing an order of magnitude higher error, confirming that fragmentation breaks bitwise determinism.
@@ 3.2 Challenge-#2: Performance Determinism
Meeting strict tail-latency SLOs is a prerequisite for ensuring satisfactory user experience in interactive deep learning workloads. However, achieving performance determinism—specifically, keeping P99 latency stable—is notoriously difficult in multi-tenant environments. While temporal sharing is the incumbent solution, it is fundamentally ill-suited for strict latency constraints. The primary limitation is that time-slicing enforces serialization: it introduces unavoidable queuing delays for applications not currently scheduled on the GPU [27, 77]. As contention increases, these accumulated wait times frequently cause tail latencies to exceed strict deadlines.
@@
  • Micro-architectural Contention: Spatial co-location implicitly multiplexes shared resources that are not explicitly partitioned, most notably Device Memory Bandwidth and Last-Level Cache capacity. As characterized in prior studies [22], a latency-sensitive inference job can suffer non-deterministic latency spikes (up to $3\times$ [43]) when co-located with a bandwidth-heavy training job.
  • Scheduler Obliviousness: Commodity GPU hardware schedulers are throughput-oriented and oblivious to high-level SLO constraints. Without explicit intervention, a high-priority kernel may suffer from head-of-line blocking behind a massive, non-preemptible background kernel in the hardware command queue [64], rendering logical priority ineffective.